# ScaleBiO: Scalable Bilevel Optimization for LLM Data Reweighting

### v1

@@ Abstract
Bilevel optimization has shown its utility across various machine learning settings, yet most algorithms in practice require second-order information, making it challenging to scale them up. Only recently, a paradigm of first-order algorithms emerged, capable of effectively addressing bilevel optimization problems. Nevertheless, the practical efficiency of this paradigm remains unverified, particularly in the context of large language models (LLMs). This paper introduces the first scalable instantiation of this paradigm called ScaleBiO, focusing on bilevel optimization for large-scale LLM data reweighting. By combining with a recently proposed memory-efficient training technique called LISA, our novel algorithm allows the paradigm to scale to 34-billion-parameter LLMs on eight A40 GPUs, marking the first successful application of bilevel optimization under practical scenarios for large-sized LLMs. Empirically, extensive experiments on data reweighting verify the effectiveness of ScaleBiO for different-scaled models, including GPT-2, LLaMA-3-8B, GPT-NeoX-20B, and Yi-34B, where bilevel optimization succeeds in filtering irrelevant data samples and selecting informative samples. Theoretically, ScaleBiO ensures the optimality of the learned data weights, along with a convergence guarantee matching the conventional first-order bilevel optimization paradigm on smooth and strongly convex objectives.
Preprint. Work in Progress.
@@ 1 Introduction
Bilevel optimization (BO) has emerged as a prominent area of research, drawing substantial attention due to its effectiveness in numerous machine learning applications, such as hyperparameter optimization (Domke, 2012; Maclaurin et al., 2015; Franceschi et al., 2017; Lorraine et al., 2020), meta-learning (Andrychowicz et al., 2016; Franceschi et al., 2018; Rajeswaran et al., 2019) and reinforcement learning (Konda and Tsitsiklis, 1999; Hong et al., 2020). In its standard formulation, bilevel optimization involves a two-level hierarchical structure, where the solution to the outer-level problem depends on the solution to the inner-level problem,
@@
Despite the inherent flexibility and applicability of bilevel optimization across a wide range of problems, its extensive utilization in large-scale problems has been relatively limited thus far. The primary obstacle hindering the scalability of bilevel optimization arises from the interdependence between the upper-level and lower-level problems. This mutual dependency introduces significant computational challenges, particularly when dealing with large-scale problem instances. The natural gradient-based iterative method of solving Problem (1) is to compute (or estimate) the hyper-gradient
$\frac{\partial\mathcal{L}(\lambda)}{\partial\lambda}=\frac{\partial L_{1}(w_{\ast}(\lambda),\lambda)}{\partial\lambda}+\frac{\partial L_{1}(w_{\ast},\lambda)}{\partial w_{\ast}}\frac{\partial w_{\ast}(\lambda)}{\partial\lambda}.$ (2)
Under certain smoothness assumptions, the derivative $\frac{\partial w_{\ast}(\lambda)}{\partial\lambda}$ exists and is derived by implicit function theorem $\frac{\partial L_{2}}{\partial w}(w_{\ast}(\lambda),\lambda)=0$ via
$\frac{\partial^{2}L_{2}(w_{\ast}(\lambda),\lambda)}{\partial^{2}w}\frac{\partial w_{\ast}(\lambda)}{\partial\lambda}+\frac{\partial^{2}L_{2}(w_{\ast}(\lambda),\lambda)}{\partial\lambda\partial w}=0.$ (3)
The main challenge lies in computing or approximating the derivative $\partial w_{\ast}(\lambda)/\partial\lambda$ in (2), where a direct computation through (3) involves the Jacobian and Hessian of the inner function $L_{2}$ and may even require the Hessian inverse. There is a line of research (Domke, 2012; Pedregosa, 2016; Grazzi et al., 2020; Lorraine et al., 2020; Franceschi et al., 2017; Shaban et al., 2019; Grazzi et al., 2020; Ghadimi and Wang, 2018; Hong et al., 2020; Yang et al., 2021; Ji et al., 2021; Chen et al., 2022) have been tempted to address this challenge. However, these works mentioned above all require the computations of Hessian, Jacobian, or their products with vectors, which can be computationally expensive and memory-intensive for large-scale problems. Recently,  Kwon et al. (2023) proposed a fully first-order method for stochastic bilevel optimization via only the first-order gradient oracle. This approach addresses the challenges associated with second-order computations and offers promising potential for stochastic bilevel optimization.
Despite these groundbreaking advancements in algorithms and theory, the practical performance of bilevel optimization in large-scale real-world settings has yet to be thoroughly investigated. Aiming to close this gap, this paper considers a practical scenario where Large Language Models (LLMs) are fine-tuned with different sources of datasets. We identify a significant challenge in determining the optimal sampling weights for each data source. For instance, Wang et al. (2024) has demonstrated that LLMs’ task-specific performance degrades in the presence of certain training datasets. However, the inclusion and combination of various datasets should intuitively enhance the models’ overall performance with proper sampling weights. This data-task misalignment poses a primary challenge in training LLMs with multiple data sources:
@@
Various methods have been proposed in attempting to address this challenge. However, they either rely on intuitive preset (Zhou et al., 2024; Muennighoff et al., 2022; Du et al., 2022a; Almazrouei et al., 2023) or lacks theoretical guarantees (Xia et al., 2024a; Xie et al., 2024; Xia et al., 2024b), leading to suboptimal sampling weights. To this end, we land bilevel optimization to the data re-weighting tasks for LLMs, aiming to overcome the limitations of existing methods. We formulate this problem into a bilevel optimization problem where the sampling weights act as the outer learnable variables. Extensive experimental results in data denoising, multi-lingual training, and instruction-following fine-tuning demonstrate the scalability and effectiveness of our approach. The primary contributions are summarized as follows:
  • We first propose the scalable instantiation of bilevel optimization on large-sized LLM training problems, which is capable of scaling to models with 34 billion parameters.
  • We successfully bridge the gap between algorithmic advancements in bilevel optimization and their application in data reweighting, allowing the optimal data weights to be learnable for large-scale LLMs.
  • We provide both experimental and theoretical results to demonstrate the efficiency of ScaleBiO. Empirically, ScaleBiO is capable of filtering irrelevant data samples and selecting informative samples. Theoretically, ScaleBiO’s convergence guarantee matches the results of Kwon et al. (2023) on smooth and strongly convex objectives.
@@ 2 Related Work
ScaleBiO (Ours) 1-st order, stochastic data reweighting Yi 34B
@@ 2.1 Bilevel Optimization
Traditional bilevel optimization algorithms are majorly categorized into two classes: 1) approximate implicit differentiable (AID) methods (Domke, 2012; Pedregosa, 2016; Grazzi et al., 2020; Lorraine et al., 2020), or 2) iterative differentiable (ITD) methods (Domke, 2012; Maclaurin et al., 2015; Franceschi et al., 2017; Shaban et al., 2019; Grazzi et al., 2020; Zhou et al., 2022a; Gao et al., 2023; Zhou et al., 2022b; Lin et al., 2023). Both approaches follow a two-loops manner and require huge computational cost for large-scale problems. To reduce the cost, attempts in stochastic bilevel optimization have been made (Ghadimi and Wang, 2018; Hong et al., 2020; Ji et al., 2021; Chen et al., 2022; Khanduri et al., 2021), which significantly improve the efficiency of traditional methods, but still lack practicality for large-scale settings due to the requirements of second-order information, such as Jacobian- and Hessian-vector products for estimating the hyper-gradient. Sow et al. (2022); Yang et al. (2023) attempt to approximate the Jacobian matrix $\nabla y^{\ast}(x)$ in (2) by finite differences, but the finite-different estimation can be sensitive to the selection of the smoothing constant and may suffer from some numerical issues in practice (Jorge and Stephen, 2006).
@@
Recently, a new paradigm of fully first-order penalty-based methods has been introduced, which reformulate the inner-level problem into the optimality constraint (Liu et al., 2022; Kwon et al., 2023; Chen et al., 2023). Liu et al. (2022) first found the hypergradient only involving first-order information, while the method only applies to deterministic functions. Kwon et al. (2023) introduced a first-order gradient-based approach that avoids the estimations of Hessian or Jacobian. This method is easily adapted and extended to stochastic bilevel optimization settings. Chen et al. (2023) provided the near-optimal sample complexity, which improves the theoretical result of (Kwon et al., 2023) in the deterministic bilevel optimization. These results verify the effectiveness of the proposed paradigm in theory, yet its practical applications in large-scale LLM settings remain unexplored.
@@
On the practical side, bilevel optimization has been explored in various NLP tasks. Somayajula et al. (2023) uses bilevel optimization to learn the task-dependent similarity structure. Although their approach demonstrates effectiveness on BERT models (Devlin et al., 2018), the finite difference approximation suffers from high error and therefore lacks the scalability in LLMs with billions of parameters. Grangier et al. (2024) adopts SOBA (Dagréou et al., 2022) to modify the training data distributions for language modeling under domain shift. However, the algorithm still requires gradient approximation and Hessian-vector products, posing challenges to scalability and engineering for large-scale problems. We summarize typical bilevel algorithms and their model sizes in Table 1, where to the best of our knowledge, no approach listed in the table has been successfully applied to over 1B-sized LLM models.
@@ 2.2 Data Reweighting
The proportion of training data sources significantly affects the performance of large language models (Du et al., 2022b; Xie et al., 2023). To this end, various methods have been proposed to reweight data sources for optimal training data mixture. For instance,  Mindermann et al. (2022) utilizes the loss gap between a trained model and a base model to identify learnable data samples, assigning them higher weights on the fly. Thakkar et al. (2023) proposes to use self-influence score to guide the reweighting in mini-batch during pretraining. Xia et al. (2024b) leverages reference losses on validation sets and adjusts the weights dynamically, adding minimal overhead to standard training. DoReMi (Xie et al., 2024) applies distributionally robust optimization (DRO) to tuning the domain weights without knowledge of downstream tasks, showing effectiveness in improving perplexity and downstream accuracy. Nevertheless, none of the aforementioned methods ensures the optimality of the learned data weights, let alone scalable experiments on over 30B-sized models.
@@ 3 Methods
In this section, we elaborate on our ScaleBiO method for finding the optimal sampling weights in training large-scaled LLMs. We first formulate this problem as a bilevel optimization problem in Section 3.1 and then develop an efficient training method for our formulation in Section 3.2.
@@ 3.1 Problem Formulation
Suppose that $m$ data sources are available for training, e.g. Alpaca (Taori et al., 2023), FLAN (Wei et al., 2021), and ShareGPT (Chiang et al., 2023), where each source $S_{i}$ is a set of $n_{i}$ examples $S_{i}=\{a_{1}^{i},a_{2}^{i},\dots,a_{n_{i}}^{i}\}$. We balance multiple training data sources by assigning each data source $S_{i}$ a sampling weight $p_{i}$ that satisfies $\sum_{i=1}^{m}p_{i}=1$.
Accordingly, each data source $S_{i}$ contributes $p_{i}|\mathcal{D}_{\text{trn}}|$ samples to the training dataset $\mathcal{D}_{\text{trn}}$. We then optimize the sampling weights to minimize the model’s loss on validation set $\mathcal{D}_{\text{val}}$. This leads to the following bilevel optimization problem:
@@
where $w$ denotes the parameters of LLM, $\{p_{i}\}$ is the probability distribution over $m$ data sources, $L_{\text{val}}$ and $L_{\text{trn}}$ respectively denote the language modeling loss on $D_{\text{val}}$ and $D_{\text{trn}}$. In addition, we introduce a trainable variable $\lambda\in\mathbb{R}^{m}$ to represent $p_{i}=e^{\lambda_{i}}/\sum_{j=1}^{m}e^{\lambda_{j}}$, so that $\sum_{i=1}^{m}p_{i}=1$ is satisfied.
@@
4:     Generating iid samples $\left\{D_{\text{tr}}^{k},D_{\text{val}}^{k}\right\}$ from training dataset $D_{\text{tr}}$ and validation dataset $D_{\text{val}}$
@@
6:     $u_{k+1}=u_{k}+U_{j_{k}}(u_{k+1}^{j_{k}}-u_{k}^{j_{k}})$
@@
8:     $w_{k+1}=w_{k}+W_{r_{k}}(w_{k+1}^{r_{k}}-w_{t}^{r_{k}})$

### v2

@@ Abstract
Bilevel optimization has shown its utility across various machine learning settings, yet most algorithms in practice require second-order information, making it challenging to scale them up. Only recently, a paradigm of first-order algorithms has emerged in the theoretical literature, capable of effectively addressing bilevel optimization problems. Nevertheless, the practical efficiency of this paradigm remains unverified, particularly in the context of large language models (LLMs). This paper introduces the first scalable instantiation of this paradigm called ScaleBiO, focusing on bilevel optimization for large-scale LLM data reweighting. By combining with a recently proposed memory-efficient training technique called LISA, our novel algorithm allows the paradigm to scale to $\sim$30B-sized LLMs on 8×8$\times$8 ×H100 GPUs, marking the first successful application of bilevel optimization under practical scenarios for large-sized LLMs. Empirically, extensive experiments on data reweighting verify the effectiveness of ScaleBiO for different-scaled models, including Llama-3-8B, Gemma-2-9B, Qwen-2-7B, and Qwen-2.5-32B, where bilevel optimization succeeds in instruction-following and math reasoning tasks, outperforming several popular baselines, including uniform sampling, influence-aware data filtering, and reference-model-based sampling methods. Theoretically, ScaleBiO ensures the optimality of the learned data weights, along with a convergence guarantee matching the conventional first-order bilevel optimization paradigm on smooth and strongly convex objectives.
ScaleBiO: Scalable Bilevel Optimization for LLM Data Reweighting
Rui Pan [^1] footnotemark: 11,   Dylan Zhang [^1] footnotemark: 11,   Hanning Zhang [^1] footnotemark: 11,   Xingyuan Pan1thanks:   Equal Contribution.1,   Minrui Xu21, Jipeng Zhang21,   Renjie Pi2,   Xiaoyu Wang21, Tong Zhang1 1University of Illinois Urbana-Champaign 2The Hong Kong University of Science and Technology {ruip4, shizhuo2, hanning5, xp12}@illinois.edu {mxubh, jzhanggr, rpi, maxywang}@ust.hk tongzhang@tongzhang-ml.org
@@ 1 Introduction
Data quality plays a crucial role in the success of Large Language Models (LLMs) (Gunasekar et al., 2023; Yang et al., 2024a; Dubey et al., 2024). Among various techniques for improving data quality, data reweighting has gained increasing attention for advancing LLMs, particularly in areas such as enhancing fairness (Roh et al., 2021, 2020, 2023), accelerating pre-training (Xia et al., 2024a; Xie et al., 2024), strengthening training robustness (Jain et al., 2024), and boosting transfer learning (Xia et al., 2024b). It is widely acknowledged that data reweighting and filtering techniques can lead to significant improvements across a diverse range of tasks (Gunasekar et al., 2023; Xu et al., 2024; Hu et al., 2024; Yang et al., 2024a; Liu et al., 2024).
On the other hand, Bilevel Optimization (BO) has emerged as a prominent area of research for solving this data re-weighting task, which draws substantial attention due to its effectiveness in numerous machine learning applications, such as hyperparameter optimization (Domke, 2012; Maclaurin et al., 2015; Franceschi et al., 2017; Lorraine et al., 2020), meta-learning (Andrychowicz et al., 2016; Franceschi et al., 2018; Rajeswaran et al., 2019) and reinforcement learning (Konda and Tsitsiklis, 1999; Hong et al., 2020). In its standard formulation, bilevel optimization involves a two-level hierarchical structure with inner-outer dependence,
@@
Despite the inherent flexibility and applicability of bilevel optimization across a wide range of problems, its extensive utilization in large-scale problems has been relatively limited thus far. The primary obstacle hindering the scalability of bilevel optimization arises from the interdependence between the upper-level and lower-level problems. The natural gradient-based iterative method of solving Problem (1) is to compute (or estimate) the hyper-gradient
$\frac{\partial\mathcal{L}(\lambda)}{\partial\lambda}=\frac{\partial L_{1}(w_{\ast}(\lambda),\lambda)}{\partial\lambda}+\frac{\partial L_{1}(w_{\ast},\lambda)}{\partial w_{\ast}}\frac{\partial w_{\ast}(\lambda)}{\partial\lambda},$ (2)
where the main challenge lies in efficiently computing or approximating the derivative $\partial w_{\ast}(\lambda)/\partial\lambda$ in (2). There is a line of research Domke (2012); Pedregosa (2016); Grazzi et al. (2020); Lorraine et al. (2020); Franceschi et al. (2017); Shaban et al. (2019); Grazzi et al. (2020); Ghadimi and Wang (2018); Hong et al. (2020); Yang et al. (2021); Ji et al. (2021); Chen et al. (2022) have been tempted to address this challenge. However, these works all require the computation of Hessian, Jacobian, or their products with vectors, which can be computationally expensive and memory-intensive for large-scale problems. Recently,  Kwon et al. (2023) proposed a fully first-order method for stochastic bilevel optimization via only the first-order gradient oracle. This approach addresses the challenges associated with second-order computations and offers promising potential for stochastic bilevel optimization.
Despite these groundbreaking advancements in algorithms and theory, the practical performance of theoretically-optimal bilevel optimization algorithms in large-scale real-world settings has yet to be thoroughly investigated. Aiming to close this gap, this paper considers a practical scenario where LLMs are fine-tuned with different sources of datasets. We identify a significant challenge in determining the optimal sampling weights for each data source. For instance, Wang et al. (2024) have demonstrated that LLMs’ task-specific performance degrades in the presence of certain training datasets. However, the inclusion and combination of various datasets should intuitively enhance the models’ overall performance with proper sampling weights. This data-task misalignment poses a primary challenge in training LLMs with multiple data sources:
@@
Various methods have been proposed in attempting to address this challenge. However, they either rely on intuitive preset (Zhou et al., 2024; Muennighoff et al., 2022; Du et al., 2022b; Almazrouei et al., 2023) or lacks theoretical guarantees (Xia et al., 2024b; Xie et al., 2024; Xia et al., 2024a), leading to suboptimal sampling weights. To this end, we test theoretical-optimal bilevel optimization in data re-weighting tasks for LLMs, aiming to overcome the limitations of existing methods. Our primary contributions are summarized as follows:
  • We propose the first scalable and theoretically-optimal instantiation of bilevel optimization on large-sized LLM training problems, which is capable of scaling to models with $\sim$30 billion parameters.
  • We successfully bridge the gap between theoretical advancements in bilevel optimization and their application in data reweighting, allowing the optimal data weights to be learnable for large-scale LLMs.
  • We provide both experimental and theoretical results to demonstrate the effectiveness of ScaleBiO. Empirically, ScaleBiO outperforms popular data filtering/reweighting baselines, including uniform sampling, LESS (Xia et al., 2024b), and RHO-LOSS (Mindermann et al., 2022a), surpassing them by a non-trivial margin of $1\%-9\%$ in GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). This superiority of ScaleBiO also holds in instruction following tasks. Theoretically, ScaleBiO’s convergence guarantee matches the results of Kwon et al. (2023) on smooth and strongly convex objectives.
@@ 2 Related Work
SAMA (Choe et al., 2023) 2-nd order, stochastic noisy fine-tuning BERT 110M
@@
(FG)2(Shen et al., 2024) 1-st order, stochastic online adaptation GPT-2-XL 1.5B
ScaleBiO (Ours) 1-st order, stochastic data reweighting Qwen-2.5-32B 32B
@@ 2.1 Bilevel Optimization
Traditional bilevel optimization algorithms are majorly categorized into two classes: 1) approximate implicit differentiable (AID) methods Domke (2012); Pedregosa (2016); Grazzi et al. (2020); Lorraine et al. (2020), or 2) iterative differentiable (ITD) methods Domke (2012); Maclaurin et al. (2015); Franceschi et al. (2017); Shaban et al. (2019); Grazzi et al. (2020). Both approaches follow a two-loops manner and require huge computational cost for large-scale problems. To reduce the cost, attempts in stochastic bilevel optimization have been made (Ghadimi and Wang, 2018; Hong et al., 2020; Ji et al., 2021; Chen et al., 2022; Khanduri et al., 2021), which significantly improve the efficiency of traditional methods, but still lack practicality for large-scale settings due to the requirements of second-order information, such as Jacobian- and Hessian-vector products for estimating the hyper-gradient. Sow et al. (2022); Yang et al. (2023) attempt to approximate the Jacobian matrix $\nabla y^{\ast}(x)$ in (2) by finite differences, but the finite-different estimation can be sensitive to the selection of the smoothing constant and may suffer from some numerical issues in practice Jorge and Stephen (2006).
@@
Recently, a new paradigm of fully first-order penalty-based methods has been introduced, which reformulate the inner-level problem into the optimality constraint Liu et al. (2022); Kwon et al. (2023); Chen et al. (2023). Liu et al. (2022) first find the hypergradient only involving first-order information, while the method only applies to deterministic functions. Kwon et al. (2023) introduced a first-order gradient-based approach that avoids the estimations of Hessian or Jacobian. This method is easily adapted and extended to stochastic bilevel optimization settings. Chen et al. (2023) provided the near-optimal sample complexity, which improves the theoretical result of Kwon et al. (2023) in the deterministic bilevel optimization. These results verify the effectiveness of the proposed paradigm in theory, yet its practical applications in large-scale LLM settings remain unexplored.
@@
On the practical side, bilevel optimization has been explored in various NLP tasks. Somayajula et al. (2023) use bilevel optimization to learn the task-dependent similarity structure. Although their approach demonstrates effectiveness on BERT models (Devlin et al., 2018), the finite difference approximation suffers from high error and therefore lacks the scalability in LLMs with billions of parameters. Grangier et al. (2024) adopt SOBA (Dagréou et al., 2022) to modify the training data distributions for language modeling under domain shift. However, the algorithm still requires gradient approximation and Hessian-vector products, posing challenges to scalability and engineering for large-scale problems. We summarize typical bilevel algorithms and their model sizes in Table 1, where to the best of our knowledge, no approach listed in the table has been successfully applied to over 3B-sized LLM models.
@@ 2.2 Data Reweighting
The proportion of training data sources significantly affects the performance of large language models (Du et al., 2022a; Xie et al., 2023). To this end, various methods have been proposed to reweight data sources for optimal training data mixture. For example,  Mindermann et al. (2022b) utilizes the loss gap between a trained model and a base model to identify learnable data samples, assigning them higher weights on the fly. Thakkar et al. (2023) propose to use a self-influence score to guide the reweighting in mini-batch during pre-training. Xia et al. (2024a) leverages reference losses on validation sets and adjusts the weights dynamically, adding minimal overhead to standard training. DoReMi (Xie et al., 2024) applies distributionally robust optimization (DRO) to tuning the domain weights without knowledge of downstream tasks. Nevertheless, none of the aforementioned methods ensures the optimality of the learned data weights, let alone scalable experiments on over 30B-sized models.
@@ 3 Methods
In this section, we elaborate on our ScaleBiO method for finding the optimal sampling weights when training large-scale LLMs. We first formulate this problem as a bilevel optimization problem in Section 3.1 and then develop an efficient training method for our formulation in Section 3.2.
@@ 3.1 Problem Formulation
Suppose that $m$ data sources are available for training, e.g. Alpaca (Taori et al., 2023), FLAN (Wei et al., 2021), and ShareGPT (Chiang et al., 2023), where each source $S_{i}$ is a set of $n_{i}$ examples $S_{i}=\{a_{1}^{i},a_{2}^{i},\dots,a_{n_{i}}^{i}\}$. The desired dataset mixture can be obtained by assigning each data source $S_{i}$ a sampling weight $p_{i}$ that satisfies $\sum_{i=1}^{m}p_{i}=1$.
Accordingly, each data source $S_{i}$ contributes $p_{i}|\mathcal{D}_{\text{trn}}|$ samples to the training dataset $\mathcal{D}_{\text{trn}}$, where the sampling weights can be optimized to minimize the model’s loss on validation set $\mathcal{D}_{\text{val}}$. This leads to the following bilevel optimization problem:
@@
where $w$ denotes the parameters of LLM, $\{p_{i}\}$ is the probability distribution over $m$ data sources, $L_{\text{val}}$ and $L_{\text{trn}}$ respectively denote the language modeling loss on $D_{\text{val}}$ and $D_{\text{trn}}$. To ensure non-negativity of the sampling weights $\{p_{i}\}_{i=1}^{m}$, an additional trainable variable $\lambda\in\mathbb{R}^{m}$ is introduced to represent $p_{i}=e^{\lambda_{i}}/\sum_{j=1}^{m}e^{\lambda_{j}}$.
@@
4:     Generating i.i.d. samples $\left\{D_{\text{tr}}^{k},D_{\text{val}}^{k}\right\}$ from training dataset $D_{\text{tr}}$ and validation dataset $D_{\text{val}}$
@@
6:     $u_{k+1}=u_{k}+U_{j_{k}}(u_{k+1}^{j_{k}}-u_{k}^{j_{k}})\text{\quad\quad\quad\quad\quad\quad\quad\quad{\color[rgb]{.5,.5,.5}$\triangleright$ Map the permuted block parameters back}}$
@@
8:     $w_{k+1}=w_{k}+W_{r_{k}}(w_{k+1}^{r_{k}}-w_{k}^{r_{k}})\text{\quad\quad\quad\quad\quad\quad\quad{\color[rgb]{.5,.5,.5}$\triangleright$ Map the permuted block parameters back}}$